Solve (x^2-11x+29)^(6x^2+x-2)=1
1
Expert's answer
2020-12-24T17:11:31-0500
"x^2-11x+29=1"

"x^2-11x+29-1=0"

"x^2-11x+28=0"

"(x-4)(x-7)=0"

"x_1=4, x_2=7"

Or


"6x^2+x-2=0, x^2-11x+29\not=0"

"(3x+2)(2x-1)=0"

"x_3=-\dfrac{2}{3}, x_4=\dfrac{1}{2}"


Let


"x^2-11x+29=-1"

"x^2-11x+29+1=0"

"x^2-11x+30=0"

"(x-5)(x-6)=0"

"x=5\ or\ x=6"

If "x=5"


"6x^2+x-2=6(5)^2+5-2=153"

"(-1)^{153}\not=1"

If "x=6"


"6x^2+x-2=6(6)^2+6-2=220"

"(-1)^{220}=1"

"x_5=6"

"\bigg\{-\dfrac{2}{3}, \dfrac{1}{2}, 4, 6,7\bigg\}"